The rate of heat conduction out of a glass window on a winter day is rapid enough to chill the air next to it. To see just how rapidly the windows transfer heat by conduction, calculate the rate of conduction in watts through a 3.89-m2 window that is 0.68 cm if the temperatures of the inner and outer surfaces are 8.71 ºC and -7.23ºC, respectively.


1
Expert's answer
2021-03-29T09:06:56-0400
"P=\frac{kA(T_h-T_c)}{d}\\
P=\frac{(0.84)(3.89)(8.71-(-7.23))}{0.0068}=7660\ W"